The present value of a perpetuity, with the first cash flow paid in 4 years time, is equivalent to receiving $100,000 in 15 years time. The perpetuity and the lump sum are of equivalent risk and have a required rate of return of 10%. What is the annual cash flow associated with the perpetuity?
1
Expert's answer
2018-01-08T09:02:07-0500
A perpetuity is a constant stream of identical cash flows with no end.
The annual cash flow of the perpetuity is CF = PV*r = 100,000*0.1 = $10,000.